# mSFT: Addressing Dataset Mixtures Overfitting Heterogeneously in Multi-task SFT

### 24 Mar 2026 v2

@@ 2 Motivation: Dataset Mixtures Overfit Heterogeneously
Multi-task SFT suffers from a fundamental misalignment between the diverse learning dynamics of individual tasks and the rigid nature of standard training paradigms. To formalize this, consider SFT of Language Models (LMs) parameterized by $\theta$ on a multi-task dataset mixture $\mathcal{D}=\bigcup_{i=1}^{N}\mathcal{D}_{i}$, which consists of NN distinct tasks. We measure training progress using a continuous compute variable $c$, generalizing training epochs into finer-grained units (e.g., fractional epochs). For any given task $i$, there exists an optimal compute $c^{*}_{i}$, defined as the stopping point where the model achieves maximum generalization on the task’s held-out test set:
@@ 3.1 Limitation of a Naïve Solution
where $w_{i}$ is the weight of the sub-dataset ii. Consequently, the optimal compute budget $c^{*}_{i}$ for any specific task $i$ is conditional on the gradient interactions from the complete mixture.
@@ 4.4 Further Analysis
Figure 5: Robustness across varying dataset sizes. $\Delta$ Accuracy of Continual SFT, IES, and mSFT relative to SFT. mSFT consistently achieves the highest performance gains across different total dataset sizes and tasks (NN), avoiding the degradation seen in Continual SFT at larger scales.
@@ F.1 Method-specific FLOPs
summing over all NN sub-datasets trained sequentially.
@@
Dynamic mixture optimization (Shin et al., 2025) via multi-armed bandits with 1-step look-ahead. At each update step (1% of total steps), the algorithm samples batches of size $B_{\text{look-ahead}}$ for all NN sub-datasets and performs forward-backward passes to estimate look-ahead rewards, incurring $8|\theta|$ FLOPs per token (2 forward pre-loss, 4 backward, 2 forward post-loss). Between updates, training proceeds with current mixture probabilities:
@@
Step 1: Identical to SRO SFT Step 1, recording per-sub-dataset peak $c_{i}^{*}$. Step 2: Rather than hard exclusions, re-trains for $C$ units with per-category sampling weight $w_{i}=c_{i}^{*}/\bar{c}$, where $\bar{c}=\frac{1}{N}\sum_{i}c_{i}^{*}$ is the mean peak across all NN sub-datasets. Early-peaking sub-datasets contribute fewer tokens; late-peaking subsets receive more exposure.

### 26 Mar 2026 v4

@@ 2 Motivation: Dataset Mixtures Overfit Heterogeneously
Multi-task SFT suffers from a fundamental misalignment between the diverse learning dynamics of individual tasks and the rigid nature of standard training paradigms. To formalize this, consider SFT of Language Models (LMs) parameterized by $\theta$ on a multi-task dataset mixture $\mathcal{D}=\bigcup_{i=1}^{N}\mathcal{D}_{i}$, which consists of $N$ distinct tasks. We measure training progress using a continuous compute variable $c$, generalizing training epochs into finer-grained units (e.g., fractional epochs). For any given task $i$, there exists an optimal compute $c^{*}_{i}$, defined as the stopping point where the model achieves maximum generalization on the task’s held-out test set:
@@ 3.1 Limitation of a Naïve Solution
where $w_{i}$ is the weight of the sub-dataset $i$. Consequently, the optimal compute budget $c^{*}_{i}$ for any specific task $i$ is conditional on the gradient interactions from the complete mixture.
@@ 4.4 Further Analysis
Figure 5: Robustness across varying dataset sizes. $\Delta$ Accuracy of Continual SFT, IES, and mSFT relative to SFT. mSFT consistently achieves the highest performance gains across different total dataset sizes and tasks ($N$), avoiding the degradation seen in Continual SFT at larger scales.
@@ F.1 Method-specific FLOPs
summing over all $N$ sub-datasets trained sequentially.
@@
Dynamic mixture optimization (Shin et al., 2025) via multi-armed bandits with 1-step look-ahead. At each update step (1% of total steps), the algorithm samples batches of size $B_{\text{look-ahead}}$ for all $N$ sub-datasets and performs forward-backward passes to estimate look-ahead rewards, incurring $8|\theta|$ FLOPs per token (2 forward pre-loss, 4 backward, 2 forward post-loss). Between updates, training proceeds with current mixture probabilities:
@@
Step 1: Identical to SRO SFT Step 1, recording per-sub-dataset peak $c_{i}^{*}$. Step 2: Rather than hard exclusions, re-trains for $C$ units with per-category sampling weight $w_{i}=c_{i}^{*}/\bar{c}$, where $\bar{c}=\frac{1}{N}\sum_{i}c_{i}^{*}$ is the mean peak across all $N$ sub-datasets. Early-peaking sub-datasets contribute fewer tokens; late-peaking subsets receive more exposure.